OpenStudy (anonymous):

what are the steps to find the answer to 5-2(x+5)=30

OpenStudy (anonymous):

you have to isolate x. so ill start you off... 5-2(x+5) = 30 3(x+5) = 30 ..what would you do next to isolate x?

OpenStudy (anonymous):

1. find the value of x which x=22 2. add the parenthesis 22+5=27 3. subtract 5-2 which equals 3 4. add 27+3=30

OpenStudy (anonymous):

umm are'nt we supposed to solve for x??

OpenStudy (anonymous):

because i would have done 3(x+5) = 30 3x +15 = 30 3x = 30-15 x = 15/3 x=3

OpenStudy (anonymous):

but it won't equal 30

OpenStudy (anonymous):

oh oh ohhh my bad. god im out of it because 15/3 is not 3 lol it's 5

OpenStudy (anonymous):

x=5 so it will equal to 30 5-2(5+5)=30 3(10) = 30 30 = 30 left side is equal to right side no??